OpenStudy (ggscupcakes443):

The dress store is having a sale where all merchandise is 1/4 off. A woman buys $24 of merchandise at sale price. What would she have had to pay at the regular price?

OpenStudy (ggscupcakes443):

@mathmate

OpenStudy (mathmate):

1/4 off means 1/4*100%=25% off. The lady paid 100-25=75% of the regular price. Use the following proportion: 75%=24 100%=x Cross multiply x=24*100/75=?

OpenStudy (ggscupcakes443):

32

OpenStudy (ggscupcakes443):

@mathmate

OpenStudy (ggscupcakes443):

32

OpenStudy (mathstudent55):

Correct

OpenStudy (ggscupcakes443):

thank you

OpenStudy (mathstudent55):

Original price = x Discounted price = 0.75x 0.75x = 24 x = 24/0.75 x = 32

OpenStudy (mathstudent55):

You're welcome.